OpenStudy (anonymous):

A container holds a combination of dimes and quarters. There are 178 coins in all, and their total value is $29.80. The system of equations models this situation, where x is the number of dimes and y is the number of quarters. x+y=178 0.1x+0.25y=29.8 How many dimes and quarters are in the container? A. 123 dimes and 55 quarters B. 114 dimes and 64 quarters C. 102 dimes and 76 quarters D. 98 dimes and 80 quarters

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@Hero

OpenStudy (mathmath333):

u have the two given equations u have to solve them \(x+y=178---(1)\\ 0.1x+0.25y=29.8\\ \dfrac{x}{10}+\dfrac{25y}{100}=\dfrac{298}{10}\\ \text{multilying both sides by 100}\\~\\ 10x+25y=2980\\~\\ \text{divide by 5}\\~\\ 2x+5y=596-----------(2)\\~\\ \)

OpenStudy (anonymous):

so how would i find the answee

OpenStudy (anonymous):

answer*

OpenStudy (mathmath333):

multilying\( (1) \)by \(-2\) \(-2x-2y=-356-------(3)\) add \((2)\) and \((3)\)

OpenStudy (anonymous):

I think it might be A or C

OpenStudy (mathmath333):

both are wrong

OpenStudy (anonymous):

B?

OpenStudy (mathmath333):

can u add (2) and (3)

OpenStudy (anonymous):

5

OpenStudy (mathmath333):

lol i m asking about the equations (2) and (3)

OpenStudy (anonymous):

im really confused can u tell me the answer and then explain i think that would be easier

OpenStudy (mathmath333):

sry i cant tell the answer , its againts the rules u need to solve these \((2x+5y)+(-2x-2y)=596+(-356)\)

OpenStudy (anonymous):

the problem is that i dont know how...

OpenStudy (mathmath333):

what is \(2-2\)

OpenStudy (anonymous):

0

OpenStudy (mathmath333):

what is \(2x-2x\)

OpenStudy (anonymous):

x

OpenStudy (mathmath333):

no its \(0\), \(x\) can be any unknown number

OpenStudy (anonymous):

ok

OpenStudy (mathmath333):

just like \(2-2\)

OpenStudy (anonymous):

ok

OpenStudy (mathmath333):

and what is \(5y-2y\)

OpenStudy (anonymous):

3y

OpenStudy (mathmath333):

good so what will be \((2x+5y)+(-2x-2y)=\)

OpenStudy (anonymous):

x-3y=249

OpenStudy (anonymous):

240*

OpenStudy (mathmath333):

\((2x+5y)+(-2x-2y)\\ =2x+5y-2x-2y\\ =3y\) is that right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i think so

OpenStudy (mathmath333):

so \(3y=240\) \(y=\dfrac{240}{3}\)

OpenStudy (anonymous):

y=80

OpenStudy (anonymous):

so that means its D?

OpenStudy (mathmath333):

ok \(x+y=178\\ x+80=178\\ x=?\)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

its D

OpenStudy (mathmath333):

u found \(x\) ?

OpenStudy (anonymous):

yes x=98

OpenStudy (mathmath333):

\(\Huge \checkmark\)

OpenStudy (anonymous):

Thanks :)